# Junk DNA Hypothesis: Pruning Small Pre-Trained Weights Irreversibly and Monotonically Impairs “Difficult" Downstream Tasks in LLMs

### v3

@@ 1 Introduction
A prevailing belief suggests that neural networks contain significant and non-significant components across their parameters that record necessary or unnecessary expertise respectively to handle a task, which forms the inherent basis of numerous model compression techniques (Han et al., 2016). Moreover, as model sizes continue to expand, the volume of redundant parameters is poised to escalate (Liu et al., 2022; Kim et al., 2021). This principle extends its effectiveness even to billion-level Large Language Models (LLMs) (Jaiswal et al., 2023b; Frantar & Alistarh, 2023; Hu et al., 2021; Lin et al., 2023b). With the negligible loss in performance stemming from the absence of non-significant components, a widely held belief has taken root: these non-significant components are essentially superfluous components that make a scant contribution to the model’s functionality. Yet we pause and pose a question: Could there be crucial facets overlooked in the context of whether these non-significant components are truly inconsequential artifacts for large-scale models?
@@ 2 Experimental Settings
Sparsity: We consider two types of sparsities: (1) Unstructured Sparsity: individual weights in the model are zeroed out independently, leading to irregular zero patterns (LeCun et al., 1990; Han et al., 2016); (2) Structured N:M Sparsity: a fine-grained sparsity pattern in which only N weights are non-zero for every continuous M weights (Nvidia, 2020; Zhou et al., 2021). We report the results of M=8 (N ranges from 7 to 1). Note that we intentionally focus on one-shot magnitude pruning, to isolate the effect of small-magnitude weights as the sole “delta" between the two models, and due to its recently observed promise to retain LLM performance (Jaiswal et al., 2023b). We additionally include two more SoTA LLM pruning methods: SparseGPT (Frantar & Alistarh, 2023), and Wanda (Sun et al., 2023) to confirm whether the validity of our hypothesis can generalize to other weight importance criteria as well.
@@ 4 Does Quantization Also Yield Monotonic Impairment Across Difficulty Spectrum?
Pruning and quantization are two popularly used remedies to reduce the overheads of billions of parameters in current LLMs. Pruning techniques primarily shrink network sizes by removing specific weights from the model – essentially setting them to zero. Quantization methods aim to quantize parameters into low bit-precision to reduce compute and memory budget. Across recently developed quantization methods for LLMs (Lin et al., 2023a; Frantar et al., 2022a; Dettmers et al., 2023), we adopted GPTQ for the quantization of all linear layers (exactly similar to our pruning settings) to study the monotonic impairment induced across the task difficulty spectrum. In this section, we firstly ask: How does the ratio of compression using quantization reflect on the performance of the downstream tasks? Does quantization strength dictate any performance pattern across the difficulty spectrum of the underlying task?
To this end, Figure 7 presents our head-to-head comparison for one-shot low-magnitude pruning and GPTQ quantization (Frantar et al., 2022b) various compression strengths. We adopted retrieval-augmented QA (task setting 3) in which task difficulty is controlled by varying the context length (token counts) accommodating the correct answer phrase, and variable option count in multiple-choice QA (task setting 2) for our experiments.
@@
The classical LTH (with iterative pruning and re-training) has not yet been studied for modern-scale LLMs due to its prohibitive computational costs. Previous work in smaller models did find reusable sparse substructures that contain sufficient information to handle a few similar tasks altogether (Chen et al., 2020; Yu et al., 2020). Recent attempt (Jaiswal et al., 2023b) demonstrated a similar property on LLMs for a range of tasks, using cheap one-shot pruning without re-training; but the authors also recognized the “sweetpoint sparsity" was correlated to the difficulty of downstream tasks involved (although they did not provide a concrete way to quantify such). As a modern LLM can perform an infinite range of downstream tasks in zero- or few-shot, Junk DNA Hypothesis can be viewed as a scaled extension of LTH, which suggests that: if the task difficulty spectrum (much broader than before) is taken into consideration, then there might not exist a single “lottery" subset of weight, that can be removed without causing irreparable damage for the performance of all tasks - although one might still be able to find a “lottery" for some (easy) tasks.
@@ A.1 Classical Pruning and Sparse Neural Networks
Pruning removes specific parts of a deep neural network, such as weights, neurons, or filters. The initial purpose of pruning is retrospectively to accelerate the model at inference time (a.k.a., post-training sparsification (Mozer & Smolensky, 1989; LeCun et al., 1990). Post-training sparsification has been well studied and results in various mature criteria that can be generally categorized into zero-order methods (Han et al., 2016; Gale et al., 2019), first-order methods (Molchanov et al., 2016; Sanh et al., 2020; Jiang et al., 2021), and second-order methods (LeCun et al., 1990; Hassibi & Stork, 1992; Dong et al., 2017) - the last usually achieve higher performance but also are more expensive due to the Hessian calculation, leading to the development of many Hessian approximation approaches (Zeng & Urtasun, 2018; Wang et al., 2019; Singh & Alistarh, 2020; Kurtic et al., 2022). The Lottery Ticket Hypothesis (LTH) (Frankle & Carbin, 2019) utilizes iterative magnitude pruning (IMP) to identify a subnetwork at initialization that can be re-trained independently to the original dense network’s performance. Sparse training (Mocanu et al., 2018; Jaiswal et al., 2022; Mostafa & Wang, 2019; Evci et al., 2020; Liu et al., 2021; Yuan et al., 2021; Yin et al., 2023; Kundu et al., 2021), on the other hand, starts with a (random) sparse network and updates network connectivity during training to search for good sparse neural network without any pre-training and dense training steps.
@@ A.2 Sparsity in Large-Scale Models
The advent of large-scale pre-trained models has led to the development of advanced post-training pruning methods, aiming to enhance the cost-effectiveness of these expansive models (Sanh et al., 2020; Chen et al., 2020; Jaiswal et al., 2023c; Zafrir et al., 2021; Kurtic et al., 2022; Xu et al., 2021; Lagunas et al., 2021; Zhang et al., 2022; Frantar et al., 2021; Jaiswal et al., 2023b; Ma et al., 2023). Among them, Frantar et al. (2021) extend second-order pruning to the BERT-level scale, enabling the pruning of blocks of weights and achieving state-of-the-art results for sparse BERT. Frantar & Alistarh (2023) introduce SparseGPT for pruning large language models (LLMs) in a single shot without requiring re-training or fine-tuning. They leverage column-wise second-order pruning, and successfully remove 100B weights from OPT-175B without a significant increase in perplexity. More recently, Sun et al. (2023) propose a straightforward pruning method that takes both weights and activations into account, demonstrating comparable performance to Frantar & Alistarh (2023). Li et al. (2022) reveal that activation sparsity is a prevalent phenomenon in Transformers (90% of intermediate output), yielding another opportunity for acceleration. Liu et al. (2023) introduce a large-scale SMC-Bench, indicating that state-of-the-art magnitude- and/or gradient-based sparse algorithms fall short when applied out-of-the-box to larger-scale models and a selected of complex downstream tasks. Our study is inspired by Liu et al. (2023), but with significantly expanded experiment scales, versatile task choices, concrete task difficulty definitions, and richer insights.

### v4

@@ 1 Introduction
A prevailing belief suggests that neural networks contain significant and non-significant components across their parameters that record necessary or unnecessary expertise respectively to handle a task, which forms the inherent basis of numerous model compression techniques (Han et al., 2016). Moreover, as model sizes continue to expand, the volume of redundant parameters is poised to escalate (Liu et al., 2022; Kim et al., 2021). This principle extends its effectiveness even to billion-level Large Language Models (LLMs) (Jaiswal et al., 2023c; Frantar and Alistarh, 2023; Hu et al., 2021; Lin et al., 2023a). With the negligible loss in performance stemming from the absence of non-significant components, a widely held belief has taken root: these non-significant components are essentially superfluous components that make a scant contribution to the model’s functionality. Yet we pause and pose a question: Could there be crucial facets overlooked in the context of whether these non-significant components are truly inconsequential artifacts for large-scale models?
@@ 2 Experimental Settings
Sparsity: We consider two types of sparsities: (1) Unstructured Sparsity: individual weights in the model are zeroed out independently, leading to irregular zero patterns (LeCun et al., 1990; Han et al., 2016); (2) Structured N:M Sparsity: a fine-grained sparsity pattern in which only N weights are non-zero for every continuous M weights (Nvidia, 2020; Zhou et al., 2021). We report the results of M=8 (N ranges from 7 to 1). Note that we intentionally focus on one-shot magnitude pruning, to isolate the effect of small-magnitude weights as the sole “delta" between the two models, and due to its recently observed promise to retain LLM performance (Jaiswal et al., 2023c). We additionally include two more SoTA LLM pruning methods: SparseGPT (Frantar and Alistarh, 2023), and Wanda (Sun et al., 2023) to confirm whether the validity of our hypothesis can generalize to other weight importance criteria as well.
@@ 4 Does Quantization Also Yield Monotonic Impairment Across Difficulty Spectrum?
Pruning and quantization are two popularly used remedies to reduce the overheads of billions of parameters in current LLMs. Pruning techniques primarily shrink network sizes by removing specific weights from the model – essentially setting them to zero. Quantization methods aim to quantize parameters into low bit-precision to reduce compute and memory budget. Across recently developed quantization methods for LLMs (Lin et al., 2023b; Frantar et al., 2022b; Dettmers et al., 2023), we adopted GPTQ for the quantization of all linear layers (exactly similar to our pruning settings) to study the monotonic impairment induced across the task difficulty spectrum. In this section, we firstly ask: How does the ratio of compression using quantization reflect on the performance of the downstream tasks? Does quantization strength dictate any performance pattern across the difficulty spectrum of the underlying task?
To this end, Figure 7 presents our head-to-head comparison for one-shot low-magnitude pruning and GPTQ quantization (Frantar et al., 2022a) various compression strengths. We adopted retrieval-augmented QA (task setting 3) in which task difficulty is controlled by varying the context length (token counts) accommodating the correct answer phrase, and variable option count in multiple-choice QA (task setting 2) for our experiments.
@@
The classical LTH (with iterative pruning and re-training) has not yet been studied for modern-scale LLMs due to its prohibitive computational costs. Previous work in smaller models did find reusable sparse substructures that contain sufficient information to handle a few similar tasks altogether (Chen et al., 2020; Yu et al., 2020). Recent attempt (Jaiswal et al., 2023c) demonstrated a similar property on LLMs for a range of tasks, using cheap one-shot pruning without re-training; but the authors also recognized the “sweetpoint sparsity" was correlated to the difficulty of downstream tasks involved (although they did not provide a concrete way to quantify such). As a modern LLM can perform an infinite range of downstream tasks in zero- or few-shot, Junk DNA Hypothesis can be viewed as a scaled extension of LTH, which suggests that: if the task difficulty spectrum (much broader than before) is taken into consideration, then there might not exist a single “lottery" subset of weight, that can be removed without causing irreparable damage for the performance of all tasks - although one might still be able to find a “lottery" for some (easy) tasks.
@@ 7 Impact Statements
Acknowledgements
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. NWO-2023.027.
@@ A.1 Classical Pruning and Sparse Neural Networks
Pruning removes specific parts of a deep neural network, such as weights, neurons, or filters. The initial purpose of pruning is retrospectively to accelerate the model at inference time (a.k.a., post-training sparsification (Mozer and Smolensky, 1989; LeCun et al., 1990). Post-training sparsification has been well studied and results in various mature criteria that can be generally categorized into zero-order methods (Han et al., 2016; Gale et al., 2019), first-order methods (Molchanov et al., 2016; Sanh et al., 2020; Jiang et al., 2021), and second-order methods (LeCun et al., 1990; Hassibi and Stork, 1992; Dong et al., 2017) - the last usually achieve higher performance but also are more expensive due to the Hessian calculation, leading to the development of many Hessian approximation approaches (Zeng and Urtasun, 2018; Wang et al., 2019; Singh and Alistarh, 2020; Kurtic et al., 2022). The Lottery Ticket Hypothesis (LTH) (Frankle and Carbin, 2019) utilizes iterative magnitude pruning (IMP) to identify a subnetwork at initialization that can be re-trained independently to the original dense network’s performance. Sparse training (Mocanu et al., 2018; Jaiswal et al., 2022; Mostafa and Wang, 2019; Evci et al., 2020; Liu et al., 2021; Yuan et al., 2021; Yin et al., 2023; Kundu et al., 2021), on the other hand, starts with a (random) sparse network and updates network connectivity during training to search for good sparse neural network without any pre-training and dense training steps.
@@ A.2 Sparsity in Large-Scale Models
The advent of large-scale pre-trained models has led to the development of advanced post-training pruning methods, aiming to enhance the cost-effectiveness of these expansive models (Sanh et al., 2020; Chen et al., 2020; Jaiswal et al., 2023b; Zafrir et al., 2021; Kurtic et al., 2022; Xu et al., 2021; Lagunas et al., 2021; Zhang et al., 2022; Frantar et al., 2021; Jaiswal et al., 2023c; Ma et al., 2023). Among them, Frantar et al. (2021) extend second-order pruning to the BERT-level scale, enabling the pruning of blocks of weights and achieving state-of-the-art results for sparse BERT. Frantar and Alistarh (2023) introduce SparseGPT for pruning large language models (LLMs) in a single shot without requiring re-training or fine-tuning. They leverage column-wise second-order pruning, and successfully remove 100B weights from OPT-175B without a significant increase in perplexity. More recently, Sun et al. (2023) propose a straightforward pruning method that takes both weights and activations into account, demonstrating comparable performance to Frantar and Alistarh (2023). Li et al. (2022) reveal that activation sparsity is a prevalent phenomenon in Transformers (90% of intermediate output), yielding another opportunity for acceleration. Liu et al. (2023) introduce a large-scale SMC-Bench, indicating that state-of-the-art magnitude- and/or gradient-based sparse algorithms fall short when applied out-of-the-box to larger-scale models and a selected of complex downstream tasks. Our study is inspired by Liu et al. (2023), but with significantly expanded experiment scales, versatile task choices, concrete task difficulty definitions, and richer insights.